OpenStudy (anonymous):

Suppose you mark n points on a circle, where n is a whole number greater than 1. The number of segments you can draw that connect these points is 1/2n^2 - 1/2n How many segments can you draw if you mark 8 points on the circle?

OpenStudy (cwrw238):

substitute n = 8 into your formula

OpenStudy (cwrw238):

hold on - the formula is a bit confusing - it looks like it should be 0.5n^2 - 0.5n in which case for 8 points there are 0.5*8^2 - 0.5*8 = 32 - 4 = 28

OpenStudy (anonymous):

yay i got the right answer

OpenStudy (cwrw238):

there way you've written it is ambiguous 1/2n could mean half * n or 1 divided by 2n

OpenStudy (anonymous):

its half

OpenStudy (cwrw238):

good